# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v1

@@ 3.3 Influence-Aware Contrastive Decoding
where $\bigvee$ represents the bitwise OR operator. The visual tokens are categorized into two groups: object-related visual tokens $\mathbf{t}^{o}$, and Unrelated Object (UO) tokens $\mathbf{t}^{u}$, defined as follows:
$\mathbf{t}^{o}=\mathbf{t}^{v}\odot\mathcal{M}_{m},\quad\mathbf{t}^{u}=\mathbf{t}^{v}\odot(1-\mathcal{M}_{m}),$ (8)
where $\odot$ indicates Hadamard product. The overall influence of these groups is defined as:
$\texttt{I}^{o}_{m}=\sum^{S}_{s=1}(\big{|}\mathbf{g}^{v}_{ms}\big{|}\odot\mathcal{M}_{ms}),\quad\texttt{I}^{u}_{m}=\sum^{S}_{s=1}(\big{|}\mathbf{g}^{v}_{ms}\big{|}\odot(1-\mathcal{M}_{ms})),$ (9)
In noun predictions, this mask removes the impact of tokens related to previously mentioned objects $\mathbf{t}^{o}$ to mitigate co-occurrence bias. For non-noun predictions, this mask calculation is not triggered; all elements in $\mathcal{M}_{m}$ are set to 0, resulting in an empty $\mathbf{t}^{o}$ set.
To jointly mitigate text-visual and co-occurrence biases, GACD adjusts the logits as follows:
$\hat{\mathcal{F}}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}=(1+\alpha_{m})\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\alpha_{m}\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m},$ (10)
where $\alpha_{m}$ is the contrastive decoding weight. This contrastive decoding aims to enlarge the Kullback-Leibler (KL) divergence  [^3] A detailed explanation of why contrastive decoding enlarges KL divergence is in supplementary Sec. B. between the contrastive logits distribution $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}$ and the original logits distribution $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}$. Here, the original logits distribution can be viewed as $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{u},\mathbf{t}^{o},\mathbf{t}^{p})_{m}$, a joint distribution with additional $\mathbf{t}^{u}$. Thus, increasing the KL divergence effectively amplifies the influence of these UO visual tokens $\mathbf{t}^{u}$.
The contrastive decoding weight, $\alpha_{m}$, controls the KL divergence distance and, from another perspective, regulates the influence of each component on the final GACD logits. The influence of UO visual tokens $\hat{\mathtt{I}}^{u}_{m}$, object-related visual tokens $\hat{\mathtt{I}}^{o}_{m}$, the prompt $\hat{\mathtt{I}}^{p}_{m}$, and previous outputs $\hat{\mathtt{I}}^{y}_{m}$ is determined by their gradient values, based on (9) and (5). Follow the chain rule, UO visual tokens are included only in the original logits $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}$ and are amplified by $(1+\alpha_{m})$, while other inputs contribute to the additional contrastive logit $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}$, experiencing smaller gradient changes.
To mitigate text-visual bias, GACD balances the influence of UO visual tokens to align with the dominant text component, defined as $\texttt{I}^{t}_{m}=\max(\texttt{I}^{p}_{m},\texttt{I}^{y}_{m})$. Aligning UO visual token influence with the question prompt $\texttt{I}^{p}_{m}$ is crucial for visually grounded responses, while balancing it with previous outputs $\texttt{I}^{y}_{m}$ is essential for preventing visual forgetting. Under this condition, $\alpha_{m}$ is calculated as:
@@
where $\tilde{\mathtt{I}}^{o}_{m}$, $\tilde{\mathtt{I}}^{p}_{m}$, and $\tilde{\mathtt{I}}^{y}_{m}$ represent the influence of object-related visual tokens, prompt, and previous output tokens, respectively, on the contrastive logits $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}$, calculated follow (5).
Furthermore, unlike existing contrastive decoding methods zhou2023analyzing ; favero2024multi ; leng2024mitigating , which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our method explicitly imposes a constraint to prevent the influence of object-related visual and prompt from becoming negative:
$\alpha_{m}=\min(\alpha_{m},\frac{\mathtt{I}^{o}_{m}}{\tilde{\mathtt{I}}^{o}_{m}-\mathtt{I}^{o}_{m}},\frac{\mathtt{I}^{p}_{m}}{\tilde{\mathtt{I}}^{p}_{m}-\mathtt{I}^{p}_{m}}).$ (12)
Additionally, to further mitigate text-visual bias in long-term generation, we propose a sample-dependent stopping criterion based on the visual influence. If the visual influence ratio $r^{v}_{m}$ for the token immediately following the end-of-sequence (EOS) token falls below a defined threshold $\epsilon$:
$r^{v}_{m}=\frac{\mathtt{I}^{v}_{m}}{\mathtt{I}^{v}_{m}+\mathtt{I}^{p}_{m}+\mathtt{I}^{y}_{m}}<\epsilon\text{ and }y_{m-1}=\text{EOS},$ (13)
@@ 4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We follow leng2024mitigating ; woo2024don ; deng2024seeing ; huang2024opera to use Amber wang2023llm , MSCOCO lin2014microsoft and LLaVa-QA90 liu2024visual datasets for the generative task, and Amber wang2023llm and POPE li2023evaluating datasets on the discriminative tasks.
@@
Implementation Details. The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling graves2013generating .
@@ 4.1 Results on Open-ended Generation
In this section, we compare against SOTA alignment-based method RLAIF-V yu2024rlaif and contrastive decoding methods VCD leng2024mitigating , M3ID favero2024multi , and AVISC woo2024don , on the AMBER and MSCOCO datasets, as presented in Tab. 1, Tab. 2. Additionally, we evaluate our method against VCD on the LLaVA-QA90 dataset, presented in Tab. 3. Our method outperforms most existing approaches across various baseline models and datasets, highlighting its robustness and reliable performance across different data types and model architectures. Specifically, we surpass image-level contrastive decoding methods like VCD and M3ID, demonstrating its effectiveness in operating at the token level and adapting to individual samples. Furthermore, compared to the token-level AVISC, our method excels, likely due to its object awareness and adaptability to fluctuating bias levels. The results further demonstrate that our method effectively mitigates hallucinations while preserving information.
@@
[…]
@@
Table 2: Results Using the CHAIR Metric on the MSCOCO Subset Following deng2024seeing .
@@
Hallucination Mitigation. Our approach achieves hallucination reductions up to $33\%$ at sentence-level (halhal in Tab. 1 and $C_{S}$ in Tab. 2) and $32\%$ at instance-level (chacha in Tab. 1 and $C_{I}$ in Tab. 2), demonstrating its effectiveness in mitigating overall hallucinations. It also effectively addresses co-occurrence hallucinations, achieving up to a $57\%$ reduction for cogcog in Tab. 1. We detail the text-visual and co-occurrence bias analysis in Sec. 5. Furthermore, an accuracy increase of up to $92\%$ in Tab. 3 underscores our model’s effectiveness in reducing hallucinations by staying closely aligned with the input image. This improvement highlights our method’s capacity to jointly mitigate multiple types of hallucinations.
@@
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following leng2024mitigating .
@@ 4.2 Results on Discriminative Task
Next, we evaluate our method on discriminative tasks using AMBER (discriminative VQA) and POPE (existence VQA) datasets, with results shown in Tab. 1 and Tab. 4. The findings confirm that our method achieves a better balance between precision and recall, resulting in consistently higher F1 scores and improved accuracy. Notably, for the high-performance baseline Intern-VL2, other methods degrade its original performance, whereas ours effectively preserves it, likely due to its bias awareness. Additionally, we observe variation in performance improvements across different question categories and baseline MLLMs.
@@
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using baseline model LLaVA-v1.5 liu2024llava . Our method improves performance across all categories, with particularly significant gains in existence, attributes, and state. These improvements can be attributed to the enhanced influence of visual tokens, which benefit categories that are more readily discernible from visual inputs.
Variation in Improvement Across MLLMs. Our method achieves the most significant improvement on mPLUG-Owl2 in object existence VQA (Tab. 4) and on LLaVA-v1.5 across various VQA categories (Tab. 1 discriminative), with minor improvement observed on InternVL2. This variation in performance correlates with the baseline visual influence ratios of the MLLMs. Fig. 3(a) presents the visual influence ratios in object existence VQA, showing that LLaVA-v1.5 exhibits the lowest visual contribution, followed by mPLUG-Owl2. This lower baseline visual influence ratio allows our method to make more impactful adjustments. In contrast, InternVL2 has an original visual influence ratio exceeding $50\%$, resulting in minimal improvement when our method is applied. The strong performance of InternVL2 can be attributed to its original high visual influence ratio, further validating the motivation behind our approach. However, this dominant visual influence ratio in InternVL2 is not consistent across other question categories or in open-ended generation tasks  [^4] Results for other question categories and visualizations of open-ended generation are in supplementary Tab. 20 and Fig. 9., which allows our method to enhance its performance on the AMBER dataset (Tab. 6).
@@ 4.3 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6 and Intern-VL2, on the Amber dataset. The results in Tab. 6 in supplementary demonstrate that our method consistently improves MLLM performance, even when their original performance is high. To further evaluate the effectiveness of our method across diverse tasks, we conduct evaluations on the MMBench liu2025mmbench and MMVet yu2023mm datasets. The results are presented in supplementary Tab. 8 and Tab. 9. The evaluation results show that our method enhances perception and recognition performance, particularly in tasks where visual contribution is straightforward. Moreover, it maintains overall task performance, demonstrating its robustness.
5 Ablation Study
In this section, we first analyze text–visual and co-occurrence biases, and then study the contributions of our proposed components. Further details on gradient computation methods, norm selection, and hyperparameter settings are provided in the supplementary material.
Figure 4: Comparison of visual influence ratios $r^{v}_{m}$ and Text-Visual GAP, with and without our GACD. (a) Without GACD, the baseline mPLUG-Owl2 shows a low initial visual influence ratio, punctuation marks and suffixes naturally have low visual influence, while objects in images initially have relatively high visual influence, which decreases as sequence length grows. Hallucinations tend to occur when the visual ratio is low. The text-visual gap confirms that text dominates the influence on predictions. (b) With GACD, the visual influence ratio increases overall and mitigates the decrease over the sequence length. The text tokens only domain influence in predictions less related to the visual, reducing hallucination.
@@ 5 Ablation Study
Co-occurrence Bias Analysis. Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’ in the image. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset lin2014microsoft . Experimental results in Fig. 5(c) indicate that when either only ‘chair’ or ‘table’ exists in the image, the hallucination rate for the other object is $23.5\%$, with a shared most influential visual token rate of $31.9\%$ in these hallucinations. This indicates that the ‘Same Most Influential Token’ phenomenon is common in co-occurrence hallucinations. Our GACD effectively reduces the hallucination where both ‘table’ or ‘chair’ are predicted in single-object images.
5.1 Component Analysis
To assess the effectiveness of each component in our proposed method, we conducted the following analyses: 1) Visual Amplification (VA) only: Applied visual amplification to all visual tokens, including during noun predictions. 2) Co-occurrence Hallucination Reduction (CR): Detected object-related visual tokens and amplified UO visual tokens during noun predictions. 3) Our full model, with early stopping (ES). Tab. 5 demonstrates that each component of our method contributes to the overall performance. VA significantly reduces hallucinations while improving object recall. CR further mitigates co-occurrence bias, a residual form of the text-visual bias addressed by VA, resulting in additional hallucination reduction. When needed, ES mechanism shortens output length to effectively reduce hallucinations with only a minor trade-off in object recall.

### v2

@@ 3.3 Influence-Aware Constrained Decoding
Object-aware Visual Token Grouping.
For each step $m$, we detect nouns in $\mathbf{y}_{<m}$ and treat each noun $y_{i}$ as an object mention. To link a mention to visual evidence, we measure the influence $I^{v}_{is}$ of visual token $s$ on step $i$, For every noun $y_{i}$, the visual token with maximal influence is selected to form a mask $\mathcal{M}_{is}$. The cumulative object mask at step $m$ aggregates all prior noun-linked tokens:
$\mathcal{M}_{ms}=\mathbf{1}\!\left[\sum_{i=1}^{m-1}\mathcal{M}_{is}>0\right],\quad\mathcal{M}_{is}=\mathbf{1}\!\left[\,y_{i}\in\text{Noun}\;\wedge\;s=\arg\max_{j}I^{v}_{ij}\,\right],$ (8)
where $\mathbf{1}[\cdot]$ is the indicator and $\wedge$ is logical AND.
The mask $\mathcal{M}_{ms}$ identifies visual tokens linked to nouns emitted before $m$. We then partition the visual tokens into object-related ($\mathbf{t}^{o}$) and unrelated-to-objects ($\mathbf{t}^{u}$) sets via a Hadamard product:
$\mathbf{t}^{o}\;=\;\mathbf{t}^{v}\odot\mathcal{M}_{m},\qquad\mathbf{t}^{u}\;=\;\mathbf{t}^{v}\odot(\mathbf{1}-\mathcal{M}_{m}).$ (9)
Object-related and unrelated influences at step $m$ are
$\texttt{I}^{o}_{m}\;=\;\sum_{s=1}^{S}\|\mathbf{g}^{v}_{ms}\|_{1}\,\mathcal{M}_{ms},\qquad\texttt{I}^{u}_{m}\;=\;\sum_{s=1}^{S}\|\mathbf{g}^{v}_{ms}\|_{1}\,(1-\mathcal{M}_{ms}).$ (10)
Masking and grouping are applied only during noun prediction (mitigate co-occurrence hallucination from object pairs). For non-noun steps, all elements in $\mathcal{M}_{m}$ are set to 0, yielding an empty $\mathbf{t}^{o}$.
Anchor-specific Influence-weighted Decoding.
Let $\mathbf{z}^{o}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{o},\mathbf{t}^{p},\mathbf{y}_{<m})$ the anchor-specific negative logits and $\mathbf{z}^{\star}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$ be the original logits. We adjust logits by
$\hat{\mathbf{z}}_{m}\;=\;(1+\alpha_{m})\,\mathbf{z}^{\star}_{m}\;-\;\alpha_{m}\,\mathbf{z}^{o}_{m},$ (11)
with $\alpha_{m}\geq 0$. In the probability space, moving along $\mathbf{z}^{\star}_{m}-\mathbf{z}^{o}_{m}$ increases the KL divergence $D_{\mathrm{KL}}(\sigma(\mathbf{z}^{\star}_{m})\|\sigma(\mathbf{z}^{o}_{m}))$ (see Sec. B). The original logits distribution $\mathbf{z}^{\star}_{m}$ can be viewed as $\pi_{\theta^{*}}(\mathbf{t}^{u},\mathbf{t}^{o},\mathbf{t}^{p},\mathbf{y}_{<m})$, i.e., a joint distribution that additionally depends on $\mathbf{t}^{u}$ compare to $\mathbf{z}^{o}_{m}$. Increasing the KL divergence therefore emphasizes the contribution of tokens $\mathbf{t}^{u}$, which are unrelated to previous mentioned objects, thereby mitigating co-occurrence bias in noun prediction. For non-noun steps, $\mathbf{t}^{u}$ coincides with $\mathbf{t}^{v}$, meaning that all visual tokens are emphasized. This adjustment helps reduce text–visual hallucination.
When analyzing token influence of $\hat{\mathbf{z}}_{m}$ in Eq. 11, the chain rule shows that $\mathbf{t}^{u}$ occur only in the original logits $\mathbf{z}^{\star}_{m}$ and are amplified by $(1+\alpha_{m})$, whereas other inputs also contribute to $\mathbf{z}^{o}_{m}$ and therefore undergo smaller influence changes. Let $\tilde{\texttt{I}}^{o}_{m},\tilde{\texttt{I}}^{p}_{m},\tilde{\texttt{I}}^{y}_{m}$ denote group influences computed on the negative branch $\mathbf{z}^{o}_{m}$ (analogous to equation 7). We then choose $\alpha_{m}$ so that the influence of $\mathbf{t}^{u}$ matches the dominant text level, $\texttt{I}^{t}_{m}\coloneqq\max(\texttt{I}^{p}_{m},\texttt{I}^{y}_{m})$. Aligning $\mathbf{t}^{u}$ influence with the question prompt $\texttt{I}^{p}_{m}$ is crucial for visually grounded responses, while balancing with previous outputs $\texttt{I}^{y}_{m}$ prevents visual forgetting.
@@
Unlike existing decoding methods Zhou et al. (2023b); Favero et al. (2024); Leng et al. (2024), which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our approach explicitly enforces non-negativity on the influence of object-related visual and prompt tokens. This corresponds to the following upper-bound condition:
$0\;\leq\;\alpha_{m}\;\leq\;\min\!\left\{\frac{\texttt{I}^{o}_{m}}{\,\tilde{\texttt{I}}^{o}_{m}-\texttt{I}^{o}_{m}\,},\;\frac{\texttt{I}^{p}_{m}}{\,\tilde{\texttt{I}}^{p}_{m}-\texttt{I}^{p}_{m}\,}\right\}.$ (13)
Sample-dependent early stopping.
Additionally, since hallucinations are more likely in long generations Zhou et al. (2023b); Peng et al. (2025), we introduce a sample-dependent stopping criterion based on visual influence. Specifically, if the visual influence ratio $r^{v}_{m}$ of the token following the end-of-sequence (EOS) falls below a threshold $\epsilon$,
$r^{v}_{m}\;\coloneqq\;\frac{\texttt{I}^{v}_{m}}{\texttt{I}^{v}_{m}+\texttt{I}^{p}_{m}+\texttt{I}^{y}_{m}}\;<\;\epsilon\quad\text{and}\quad y_{m-1}=\text{EOS}.$ (14)
@@ 4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We use Amber Wang et al. (2023), MSCOCO Lin et al. (2014) and LLaVa-QA90 Liu et al. (2024b) datasets for the generative task, and Amber Wang et al. (2023) and POPE Li et al. (2023b) datasets on the discriminative tasks.
@@
Implementation Details. The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling Graves (2013).
@@ 4.1 Results on Open-ended Generation
In this section, we compare against SOTA alignment-based method RLAIF-V and contrastive decoding methods VCD, M3ID, and AVISC, on the AMBER and MSCOCO datasets, as presented in Tab. 1, Tab. 2. Additionally, we evaluate our method against VCD on the LLaVA-QA90 dataset, presented in Tab. 3. Our method outperforms most existing approaches across various baseline models and datasets, highlighting its robustness and reliable performance across different data types and model architectures. Specifically, we surpass image-level contrastive decoding methods like VCD and M3ID, demonstrating its effectiveness in operating at the token level and adapting to individual samples. Furthermore, compared to the token-level AVISC, our method excels, likely due to its object awareness and adaptability to fluctuating bias levels. The results further demonstrate that our method effectively mitigates hallucinations while preserving information.
@@
cha $\downarrow$ cov $\uparrow$ hal $\downarrow$ cog $\downarrow$ acc $\uparrow$ $\uparrow$ $\uparrow$ F1 $\uparrow$
@@
Table 2: Results Using the CHAIR Metric on the MSCOCO Subset Following Deng et al. (2024).
@@
Hallucination Mitigation. Our approach reduces hallucination by up to $33\%$ at sentence-level (halhal in Tab. 1 and $C_{S}$ in Tab. 2) and $32\%$ at instance-level (chacha in Tab. 1 and $C_{I}$ in Tab. 2), demonstrating its effectiveness in mitigating overall hallucinations. It also effectively mitigates co-occurrence hallucinations, with reductions of up to $57\%$ for cogcog in Tab. 1. Accuracy gains of up to $92\%$ (Tab. 3) further demonstrate that our model improves alignment with the input image, highlighting its ability to jointly address text–visual and co-occurrence biases (Sec. 4.4).
@@
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following Leng et al. (2024).
@@ 4.2 Results on Discriminative Task
We next evaluate our method on discriminative tasks using AMBER (discriminative VQA) and POPE (existence VQA), with results shown in Tab. 1 and Tab. 4. Our approach achieves a better balance between precision and recall, yielding consistently higher F1 scores and improved accuracy. Notably, unlike competing methods that degrade Intern-VL2, ours preserves its performance via bias awareness, though improvements differ across categories and MLLMs.
@@
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using LLaVA-v1.5 Liu et al. (2024a). Our method improves performance across all categories, with the largest gains in existence, attributes, and state—categories strongly tied to visual cues, benefiting from enhanced visual token influence.
Variation in Improvement Across MLLMs. Our method achieves the most significant improvement on mPLUG-Owl2 in object existence VQA (Tab. 4) and on LLaVA-v1.5 across various VQA categories (Tab. 1 discriminative), with minor improvement observed on InternVL2. This variation in performance correlates with the baseline visual influence ratios of the MLLMs. Fig. 3(a) presents the visual influence ratios in object existence VQA, showing that LLaVA-v1.5 exhibits the lowest visual contribution, followed by mPLUG-Owl2. This lower baseline visual influence ratio allows our method to make more impactful adjustments. In contrast, InternVL2 has an original visual influence ratio exceeding $50\%$, resulting in minimal improvement when our method is applied. The strong performance of InternVL2 can be attributed to its original high visual influence ratio, further validating the motivation behind our approach. However, this dominant visual influence ratio in InternVL2 is not consistent across other question categories or in open-ended generation tasks  [^3] Results for other question categories and visualizations of open-ended generation are in supplementary., which allows our method to enhance its performance on the AMBER dataset (Tab. 6).
@@ 4.3 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6, Qwen2-VL and Intern-VL2, on the Amber dataset. As shown in Tab. 6 (supplementary), our method consistently improves performance, even for strong baselines. To test generality, we also evaluate on MMBench Liu et al. (2025), MMVet Yu et al. (2023), ScienceQA Lu et al. (2022), and VizWiz Gurari et al. (2018). Results (supplementary) show that our approach enhances perception and recognition, especially for tasks with clear visual grounding, while preserving overall task performance and demonstrating robustness.
4.4 Ablation Study
In this section, we first analyze text–visual and co‐occurrence biases, and then study the contributions of our proposed components. Further details on gradient computation methods, norm selection, and hyperparameter settings are provided in the supplementary material.
Figure 4: Comparison of visual influence ratios $r^{v}_{m}$ and Text-Visual GAP, with and without our GACD. (a) Without GACD, mPLUG-Owl2 shows a low initial visual influence ratio, punctuation marks and suffixes naturally have low visual influence, while objects start with higher influence that declines as the sequence grows. Hallucinations tend to occur when the visual ratio is low. The text-visual gap confirms that text dominates the influence on predictions. (b) With GACD, the visual influence ratio increases overall and mitigates the decrease over the sequence length. The text tokens only domain influence in predictions less related to the visual, reducing hallucination.
@@ 4.4 Ablation Study
Co-occurrence Bias Analysis. Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset Lin et al. (2014). Results in Fig. 5(c) indicate that when only ‘chair’ or ‘table’ exists in the image, the hallucination rate for the other object is $23.5\%$, with a $31.9\%$ rate of sharing the same most influential token. This indicates that the ‘Same Most Influential Token’ phenomenon is common in co-occurrence hallucinations. Our GACD effectively reduces the hallucination where both ‘table’ or ‘chair’ are predicted in single-object images.
Component Analysis. To assess the effectiveness of each component in our proposed method, we conducted the following variants: 1) Visual Amplification (VA) only: visual amplification is applied to all visual tokens ($\mathbf{t}^{v}$), including during noun predictions. 2) Co-occurrence Hallucination Reduction (CR): object-related visual tokens are detected, and $\mathbf{t}^{u}$ is amplified during noun predictions. 3) Our full model, with early stopping (ES). Tab. 5 demonstrates that each component of our method contributes to the overall performance. VA significantly reduces hallucinations while improving object recall. CR further mitigates co-occurrence bias, a residual form of the text-visual bias addressed by VA, resulting in additional hallucination reduction. Both VA and CR achieve these gains without introducing trade-offs. When necessary, the ES mechanism shortens outputs to effectively reduce hallucinations, with only a slight recall trade-off that remains acceptable for most MLLMs.